The length of a rectangle is twice its width. If the perimeter of the rectangle is 60 ft , find its area
Answer by drj(1380)

The length of a rectangle is twice its width. If the perimeter of the rectangle is 60 ft, find its area

Step 1. The perimeter P means adding up all the four sides of a rectangle.

Step 2. Let w be the width.

Step 3. Let 2w be the length

Step 4. Then, P=w+w+2w+2w=60.

Step 5. Solving yields the following steps

Solved by pluggable solver: EXPLAIN simplification of an expression
Your Result:
Simplify w+w+2w+2w=60
Your Answer:
  • Solutions: w=10.
  • Graphical form: Equation was fully solved.
  • Text form: w+w+2w+2w=60 simplifies to 0=0
  • Cartoon (animation) form:
    simplify_cartoon( w+w+2w+2w=60 )

Detailed explanation:


Look at .
Eliminated similar terms ,,, replacing them with
It becomes .

Look at .
Added fractions or integers together
It becomes .

Look at .
Remove unneeded parentheses around factor
It becomes .

Look at .
Moved these terms to the left
It becomes .

Look at .
Solved linear equation equivalent to 6*w-60 =0
It becomes .
Result:
Solutions: w=10.

Done!



For w=10 , then 2w=20 and P=2(10+20)=60 which is a true statement.

Step 6. Area A is given as A=length * width = 20*10=200 square feet.

Step 7. ANSWER: The area is 200 square feet.

I hope the above steps were helpful.

For free Step-By-Step Videos on Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra or for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

And good luck in your studies!

Respectfully,
Dr J


http://www.FreedomUniversity.TV
http://www.Facebook.com/FreedomUniversity.TV
http://www.Twitter.com/FreedomUTV